### v2

# Decentralized projected Riemannian gradient methodKangkang Deng, Jiang Hu

Abstract
This paper studies decentralized optimization over a compact submanifold within a communication network of $n$ nodes, where each node possesses a smooth non-convex local cost function, and the goal is to jointly minimize the sum of these local costs. We focus particularly on the online setting, where local data is processed in real-time as it streams in, without the need for full data storage. We propose a decentralized projected Riemannian stochastic recursive momentum (DPRSRM) method that employs local hybrid stochastic gradient estimators and uses the network to track the global gradient. DPRSRM achieves an oracle complexity of $\mathcal{O}(\epsilon^{-\frac{3}{2}})$, outperforming existing methods that have at most $\mathcal{O}(\epsilon^{-2})$ complexity. Our method requires only $\mathcal{O}(1)$ gradient evaluations per iteration for each local node and does not require restarting with a large batch gradient. Furthermore, we demonstrate the effectiveness of our proposed methods compared to state-of-the-art ones through numerical experiments on principal component analysis problems and low-rank matrix completion.
keywords:
decentralized optimization, Riemannian manifold, gradient tracking, consensus
{AMS}
90C06, 90C22, 90C26, 90C56
1 Introduction
Decentralized optimization has emerged as a prominent area of research, particularly for its application in large-scale systems, such as sensor networks, distributed computing, and machine learning. In these contexts, data is often partitioned across numerous nodes, rendering centralized optimization approaches impractical due to challenges such as privacy limitations and restricted computational resources. In this work, we are concerned with the distributed smooth optimization on a compact submanifold
(1.1) $\min$ $\frac{1}{n}\sum_{i=1}^{n}f_{i}(x_{i}),$
$\mathrm{s.t.}$ $x_{1}=\cdots=x_{n},\;\;$
$x_{i}\in\mathcal{M},\;\;\forall~{}i=1,2,\ldots,n,$
where $n$ is the number of nodes, $f_{i}$ is the smooth nonconvex local objective at the $i$-th node, and $\mathcal{M}$ is a (nonconvex) compact smooth submanifold embedded in $\mathbb{R}^{d\times r}$ with the extrinsic dimensions $(d,r)$, e.g., the Stiefel manifold ${\rm St}(d,r):=\{x\in\mathbb{R}^{d\times r}:x^{\top}x=I_{r}\}$. Problem (1.1) is prevalent in machine learning, signal processing, and deep learning, see, e.g., the principal component analysis [37], the low-rank matrix completion [4, 22, 12], the low-dimensional subspace learning [1, 24], and the deep neural networks with batch normalization [8, 19]. One challenge in solving (1.1) comes from the nonconvexity of the manifold constraint, which causes difficulty in achieving the consensus [7, 12, 21, 13].
In this paper, we investigate an online setting where each node $i$ interacts with its local cost function $f_{i}$ through a stochastic first-order oracle (SFO). This SFO setting is particularly relevant in various online learning and expected risk minimization problems, where the noise introduced by the SFO stems from the variability of sampling over streaming data received at each node. A notable example is online principal component analysis [5]. Our primary focus is on the oracle complexity, defined as the total number of SFO queries required at each node to compute an $\epsilon$-stationary point tuple $\{x_{1},\cdots,x_{n}\}$, as formalized in Definition 2.1.
1.1 Related works
Decentralized optimization in Euclidean space (i.e., $\mathcal{M}=\mathbb{R}^{d\times r}$) has been extensively studied over the past few decades (see, e.g., [3, 29, 36, 25, 15, 32, 33, 38, 18, 39, 27, 30]). However, since problem (1.1) involves a manifold $\mathcal{M}$, which is often nonlinear and nonconvex, these works may fail when directly applied to solve problem (1.1).
Perhaps the earliest works for solving (1.1) are [24, 28]. However, these methods require an asymptotically infinite number of consensus steps for convergence, which limits their practical applicability. For the case where $\mathcal{M}$ is the Stiefel manifold, [7] propose a decentralized Riemannian gradient descent method and its gradient-tracking version. To use a single step of consensus, augmented Lagrangian methods [34, 35] are also investigated, where a different stationarity is used. [31] propose a decentralized retraction-free gradient tracking algorithm, and show that it exhibits ergodic $\mathcal{O}(1/K)$ convergence rate. However, these studies rely on the orthogonal structure of the Stiefel manifold. Recently, [12] used the projection operators instead of retractions and expanded the distributed Riemannian gradient descent algorithm and the gradient tracking version to the compact submanifolds of Euclidean space. Moreover, the integration of decentralized manifold optimization with other algorithms has also been proposed, including the conjugate gradient algorithm [6] and the natural gradient method [21]. Furthermore, [20] achieves single-step consensus for the general compact submanifold by carefully elaborating on the smoothness structure and the asymptotic 1-Lipschitz continuity of the projection operator associated with the submanifold geometry.
Several studies have focused on the finite-sum setting of problem (1.1), where $f_{i}=\frac{1}{m}\sum_{r=1}^{m_{i}}f_{i,r}$. [7] propose a decentralized Riemannian stochastic gradient descent method. By combining the variable sample size gradient approximation method with the gradient tracking dynamic, [40] propose a distributed Riemannian stochastic optimization algorithm on the Stiefel manifold. Although both methods can also be used in the online setting, the oracle complexities are $\mathcal{O}(\epsilon^{-2})$, which is not optimal. It is worth noting that the decentralized variance reduced method [35] has been studied. However, they need to periodically calculate the full gradient, which is not suitable for the online setting.
1.2 Contribution
In this paper, we propose DPRSRM, a novel online variance-reduced method for decentralized non-convex manifold optimization with stochastic first-order oracles (SFO).
  • To achieve fast and robust performance, the DPRSRM algorithm is built upon gradient tracking [7, 12] and a stochastic gradient momentum estimator [10, 17], which can be viewed as online variance reduction method. The only existing decentralized stochastic variance-reduced manifold optimization algorithm is the VRSGT proposed by [35]. Note that VRSGT is a double-loop algorithm that requires very large minibatch sizes. Conversely, the proposed DPRSRM is a single-loop algorithm with $\mathcal{O}(1)$ oracle queries per update. Numerical experiments demonstrate the effectiveness of the proposed methods compared to state-of-the-art ones through eigenvalue problems and low-rank matrix completion.
  • Our algorithm achieves an oracle complexity of $\mathcal{O}(\epsilon^{-\frac{3}{2}})$. A comparison of the oracle complexity of DPRSRM with related algorithms is provided in Table 1, where DPRSRM achieves a lower oracle complexity than the existing decentralized stochastic manifold optimization algorithms [12, 7, 35, 40]. Moreover, DPRSRM uses a single step of consensus to achieve communication, compared to other project/retraction algorithms [7, 12, 40] that need $\log_{\sigma_{2}}(\frac{1}{2\sqrt{n}})$ rounds of consensus, where $\sigma_{2}$ is the second largest singular value of the communication graph matrix.
Algorithm Manifold types Communication Tracking VR Oracle
[7] Stiefel manifold multiple $\mathcal{O}(\epsilon^{-2})$
[40] Stiefel manifold multiple $\mathcal{O}(\epsilon^{-2})$
This paper compact submanifold single $\mathcal{O}(\epsilon^{-3/2})$
Table 1: Comparison of the oracle complexity results of Riemannian online decentralized methods. “Communication” means rounds of communications per iteration. “Tracking” denotes the gradient tracking, “VR” denotes variance reduction. We do not list the work in [35] since they focus on the finite-sum setting and are not applicable to the online setting.
1.3 Notation
[…]
2 Preliminary
This section introduces the definition of stationary point for problem (1.1), and gives a key property for compact submanifolds, i.e., proximal smoothness.
2.1 Stationary point
Let $x_{1},\cdots,x_{n}\in\mathcal{M}$ represent the local copies of each node. Let $\mathcal{P}_{\mathcal{M}}$ be the orthogonal projection onto $\mathcal{M}$. Note that for $\{x_{i}\}_{i=1}^{n}\subset\mathcal{M}$,
${\rm argmin}_{y\in\mathcal{M}}\sum_{i=1}^{n}\|y-x_{i}\|^{2}=\mathcal{P}_{\mathcal{M}}(\hat{x}).$
Any element $\bar{x}$ in $\mathcal{P}_{\mathcal{M}}(\hat{x})$ is the induced arithmetic mean of $\{x_{i}\}_{i=1}^{n}$ on $\mathcal{M}$ [26]. Let $f(z):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(z)$. The $\epsilon$-stationary point of problem (1.1) is defined as follows.
Definition 2.1 ([12]).
The set of points $\{x_{1},x_{2},\cdots,x_{n}\}\subset\mathcal{M}$ is called an $\epsilon$-stationary point of (1.1) if there exists a $\bar{x}\in\mathcal{P}_{\mathcal{M}}(\hat{x})$ such that
$\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\|x_{i}-\bar{x}\|^{2}\right]\leq\epsilon\quad{\rm and}\quad\mathbb{E}[\|\mathrm{grad}f(\bar{x})\|^{2}]\leq\epsilon.$
We refer to these two terms as consensus error and optimality error, respectively.
2.2 Proximal smoothness
Proximal smoothness is an effective tool for addressing the nonconvex nature of manifold constraints within decentralized optimization settings [12]. Define the distance from a point $x\in\mathbb{R}^{d\times r}$ to the manifold $\mathcal{M}$ and the nearest-point projection of $x$ onto $\mathcal{M}$ as $\text{dist}(x,\mathcal{M}):=\inf_{y\in\mathcal{M}}\|y-x\|\;\;\text{and}\;\;\mathcal{P}_{\mathcal{M}}(x):=\arg\min_{y\in\mathcal{M}}\|y-x\|,$ respectively. For a given number $R>0$, the $R$-tube around $\mathcal{M}$ is defined as the set $U_{\mathcal{M}}(R):=\{x:\text{dist}(x,\mathcal{M})<R\}.$
A closed set $\mathcal{M}$ is said to be $R$-proximally smooth if the projection $\mathcal{P}_{\mathcal{M}}(x)$ is unique whenever $\text{dist}(x,\mathcal{M})<R$. Following [9], an $R$-proximally smooth set $\mathcal{M}$ satisfies that for any real ${\delta}\in(0,R)$,
(2.1) $\left\|\mathcal{P}_{\mathcal{M}}(x)-\mathcal{P}_{\mathcal{M}}(y)\right\|\leq\frac{R}{R-{\delta}}\|x-y\|,~{}~{}\forall x,y\in\bar{U}_{\mathcal{M}}({\delta}),$
where $\bar{U}_{\mathcal{M}}(\delta):=\{x:\text{dist}(x,\mathcal{M})\leq\delta\}$. For instance, the Stiefel manifold is known to be a 1-proximally smooth set [2], which also provides additional examples of rank manifolds along with their specific proximal smoothness radii.
3 Problem setup and the proposed DPRSRM
In this section, we present the problem setup considered in this paper, outlining the assumptions for the objective function and the communication network. Building on this setup, we then develop a decentralized algorithm to solve problem (1.1) and provide the main convergence rate results.
3.1 Problem setup
Let us start with assumptions on problem (1.1), based on the fact that any compact $C^{2}$-submanifolds in Euclidean space belong to proximally smooth set [2, 9, 11].
Assumption 3.1.
Assume that $\mathcal{M}$ is proximally smooth with radius $R$. Each objective function $f_{i}$ is gradient Lipschitz continuous with modulus $L_{f}$ on the convex hull of $\mathcal{M}$, denoted by ${\rm conv}(\mathcal{M})$. Moreover, the objective function $f(\mathbf{x})$ has an optimal value $f_{*}$ over $\mathcal{M}^{n}$.
Under this assumption, the following Riemannian quadratic upper bound for $f_{i}$ has been estabilshed in [12].
Lemma 3.2 ([12], Lemma 2).
Under Assumption 3.1, there exists $L_{g}$, for any $x,y\in\mathcal{M}$, the following inequality holds:
(3.1) $f_{i}(y)-f_{i}(x)\leq\left<\mathrm{grad}f_{i}(x),y-x\right>+\frac{L_{g}}{2}\|y-x\|^{2},~{}i\in[n].$
Moreover, there exists a constant $L_{G}>0$ such that
(3.2) $\|\mathrm{grad}f_{i}(x)-\mathrm{grad}f_{i}(y)\|\leq L_{G}\|x-y\|,~{}i\in[n].$
We now present the assumption for the communication network. Denote by the undirected graph $G:=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}=\{1,2,\ldots,n\}$ is the set of all nodes and $\mathcal{E}$ is the set of edges. Let $W$ be the mixing matrix associated with $G$ and $W_{ij}\neq 0$ if there is an edge $(i,j)\in\mathcal{E}$ and $W_{ij}=0$ otherwise. We use the following standard assumptions on $W$, see, e.g., [7, 39].
Assumption 3.3.
We assume that the undirected graph $G$ is connected and $W$ is doubly stochastic, i.e., (i) $W=W^{\top}$; (ii) $W_{ij}\geq 0$ and $01>W_{ii}>$ 0 for all $i,j$; (iii) Eigenvalues of $W$ lie in $(-1,1]$. In addition, the second largest singular value $\sigma_{2}$ of $W$ satisfies in $\sigma_{2}\in[0,1)$.
Consider a sufficiently rich probability space $\{\Omega,\mathbb{P},\mathcal{F}\}$. For a given decentralized algorithm, we assume it generates an iterative sequence $\{x_{i,k}\}_{k\geq 0}$, where $x_{i,k}$ denotes the $k$-th iteration at node $i$. At each step, node $i$ observes a random vector $\xi_{i,k}$. We then define an increasing sequence of sub-$\sigma$-algebras within $\mathcal{F}$, constructed from the random vectors observed in succession by the network nodes: for all $k\geq 1$, $\mathcal{F}_{0}:=\{\Omega,\emptyset\},\quad\mathcal{F}_{k}:=\sigma\left(\{\xi_{i,0},\xi_{i,1},\ldots,\xi_{i,k-1}:i\in\mathcal{V}\}\right),$ where $\emptyset$ represents the empty set. The following assumptions are made regarding the stochastic gradient $\nabla f_{i}(x,\xi_{i,k})$:
Assumption 3.4.
For any $\mathcal{F}_{k}$–measurable variable $x\in\mathcal{M}$ and $k\geq 1$, the algorithm generates a sample $\xi_{i,k}\sim\Omega$ for each node $i$ and returns a stochastic gradient $\mathrm{grad}f_{i}(x,\xi_{i,k})$, there exists a parameter $\nu=\frac{1}{n}\sum_{i=1}^{n}\nu_{i}^{2}$ such that
(3.3) $\mathbb{E}\left[\mathrm{grad}f_{i}(x,\xi_{i,k})|\mathcal{F}_{k}\right]=\mathrm{grad}f_{i}(x),$
(3.4) $\mathbb{E}\left[\|\mathrm{grad}f_{i}(x,\xi_{i,k})-\mathrm{grad}f_{i}(x)\|^{2}|\mathcal{F}_{k}\right]\leq\nu_{i}^{2}.$
Moreover, the collection $\{\xi_{i,k}:i\in\mathcal{V},k\geq 1\}$ consists of independent random variables and $\mathrm{grad}f_{i}(x,\xi_{i,k})$ is the mean-squared $\bar{L}$-Lipschitz:
(3.5) $\mathbb{E}[\|\mathrm{grad}f_{i}(x,\xi_{i,k})-\mathrm{grad}f_{i}(y,\xi_{i,k})\||\mathcal{F}_{k}]\leq\bar{L}\mathbb{E}[\|x-y\|].$
To measure the oracle complexity, we give the definition of a stochastic first-order oracle (SFO) for (1.1).
Definition 3.5 (stochastic first-order oracle).
For the problem (1.1), a stochastic first-order oracle for each node $i$ is defined as follows: compute the Riemannian gradient $\mathrm{grad}f_{i}(x,\xi_{i})$ given a sample $\xi_{i}\in\Omega$.
3.2 The Algorithm
In this subsection, we introduce the Decentralized Projected Riemannian Stochastic Recursive Momentum (DPRSRM) method for addressing (1.1), along with its associated convergence results. Inspired by the notable effectiveness of variance reduction and gradient tracking in decentralized frameworks [35, 12, 7], we seek a novel combination of variance reduction and gradient tracking for decentralized online problems on compact submanifolds for improving the oracle complexity. In particular, we focus on the following stochastic gradient estimator using the momentum technique introduced in [10, 17, 14]:
(3.6) $q_{i,k}=$ $\mathrm{grad}f_{i}(x_{i,k},\xi_{i,k})$
$+(1-\tau)(d_{i,k-1}-\mathrm{grad}f_{i}(x_{i,k-1},\xi_{i,k})).$
We note that (3.6) can be rewritten as
(3.7) $q_{i,k}=$ $\tau\mathrm{grad}f_{i}(x_{i,k},\xi_{i,k})+(1-\tau)(d_{i,k-1}$
$+\mathrm{grad}f_{i}(x_{i,k},\xi_{i,k})-\mathrm{grad}f_{i}(x_{i,k-1},\xi_{i,k})),$
which hybrids stochastic gradient $\mathrm{grad}f_{i}(x_{i,k},\xi_{i,k})$ with the recursive gradient estimator in RSARAH/SRG [16, 41] for $\tau\in[0,1]$. Since the direction $q_{i,k}$ may be unbounded, we introduce a clipped gradient estimator $d_{i,k}$
(3.8) $d_{i,k}=\left\{\begin{array}[]{cc}q_{i,k}&\text{if}~{}\|q_{i,k}\|\leq B,\\ B\frac{q_{i,k}}{\|q_{i,k}\|}&\text{otherwise},\end{array}\right.$
where $B>0$ is a user-defined constant. To further reduce the variance of the gradient estimator in different nodes, we compute the gradient tracking iteration as follows:
(3.9) $s_{i,k}=\sum_{j=1}^{n}W_{ij}s_{j,k-1}+d_{i,k}-d_{i,k-1},~{}i\in[n].$
A crucial advantage of gradient tracking-type methods lies in the applicability of the use of a constant step size $\alpha$. Since $s_{i,k}$ may not remain in the tangent space $T_{x_{i,k}}\mathcal{M}$, we introduce its tangent space projection $v_{i,k}=\mathcal{P}_{T_{x_{i,k}}(s_{i,k})}$ and update the new iterate $x_{i,k+1}$ as follows:
(3.10) $x_{i,k+1}=\mathcal{P}_{\mathcal{M}}(\sum_{j=1}^{n}W_{ij}x_{j,k}-\alpha v_{i,k}),~{}i\in[n],$
where the projection-based average [12, 20], i.e., $\mathcal{P}_{\mathcal{M}}(\sum_{j=1}^{n}W_{ij}x_{j,k})$, is adopted to ensure the decrease in the consensus error among nodes. For ease of analysis, we stack variables in each node and rewrite (3.6), (3.9) and (3.10) as
[…]
where $\mathcal{P}_{T_{\mathbf{x}}\mathcal{M}^{n}}:=\mathcal{P}_{T_{x_{1,k}}\mathcal{M}}\times\dots\times\mathcal{P}_{T_{x_{n,k}}\mathcal{M}}$ and $\mathbf{\xi}_{k}=\{\xi_{i,k}\}_{i\in\mathcal{V}}$. The overall algorithm is given in Algorithm 1.
Algorithm 1 The DPRSRM for solving (1.1)
0:  Initial point $\mathbf{x}_{0}=\bar{\mathbf{x}}_{0}\in\mathcal{M}$, $\mathrm{\bf s}_{-1}=\mathbf{d}_{-1}=\mathbf{0}$, $\alpha,\tau$.
1:  Sample $\xi_{i,0}$, let $s_{i,0}=d_{i,0}=\mathrm{grad}f_{i}(x_{i,0},\xi_{i,0})$.
2:  $v_{i,0}=\mathcal{P}_{T_{x_{i,0}}\mathcal{M}}(s_{i,0})$.
3:  $x_{i,1}=\mathcal{P}_{\mathcal{M}}(\sum_{j=1}^{n}W_{ij}x_{j,0}-\alpha v_{i,0})$.
4:  for $k=1,2,\cdots,K$ do
5:     Update stochastic gradient estimator $q_{i,k}$ via (3.6)
6:     Update the clipped gradient estimator $d_{i,k}$ via (3.8).
7:     Update Riemannian gradient tracking $s_{i,k}$ via (3.9).
8:     Project onto tangent space: $v_{i,k}=\mathcal{P}_{T_{x_{i,k}}\mathcal{M}}(s_{i,k})$.
9:     Update new iterate $x_{i,k+1}$ via (3.10).
10:  end for
3.3 Main results
The convergence analysis of Algorithm 1 can be divided into two parts: the consensus error and the optimality error, as defined in Definition 2.1. Let us first focus on the consensus error. We define a neighborhood around $\mathbf{x}\in\mathcal{M}^{n}$ as follows:
(3.12) $\mathcal{N}(\delta):=\{\mathbf{x}\in\mathcal{M}^{n}:\|\mathbf{x}-\bar{\mathbf{x}}\|\leq\delta\}.$
Note that if $\mathbf{x}\in\mathcal{N}(\delta)$, then $x_{i}\in\bar{U}_{\mathcal{M}}(\delta)$ for any $i\in[n]$.
By appropriately selecting the step size $\alpha$ and an integer $t$, and initializing with $\mathbf{x}_{0}\in\mathcal{N}(\delta)$, we can establish an explicit relationship between the consensus error and the step size. The proof is given in Section 4.1.
Theorem 3.6 (Consensus error).
Let $\{\mathbf{x}_{k}\}_{k}$ be the sequence generated by Algorithm 1. Suppose that Assumptions 3.1 and 3.3 hold. Define $\rho:=\frac{R}{R-2\delta}\sigma_{2}$, $D:=\frac{(3L)^{2}}{(1-\sigma_{2})^{2}}$ and $C:=\frac{D}{(1-\rho)^{2}\sigma_{2}^{2}}$. Let $\alpha$ and $\delta$ satisfy that
(3.13) $\alpha$ $\leq\min\left\{\frac{1}{4L},\frac{\delta}{\sqrt{nD}},\frac{(R(1-\sigma_{2})-2\delta)\delta}{R\sqrt{nD}}\right\},$
$\delta$ $<\frac{1}{2}\min\left\{R,R(1-\sigma_{2})\right\}.$
If $\mathbf{x}_{0}\in\mathcal{N}(\delta)$, it holds that
(3.14) $\frac{1}{n}\|\bar{\mathbf{x}}_{k}-\mathbf{x}_{k}\|^{2}\leq C\alpha^{2}.$
Now we present the following main theorem of the DPRSRM. The proof is given in Section 4.2. For the ease of analysis, we define
(3.15) $L:=\max\{L_{g},L_{G},\bar{L}c\},$
where $L_{g},L_{G}$ are given in Lemma 3.2 and $\bar{L}$ occurred in Assumption 3.4.
Theorem 3.7 (Optimality error).
Let $\{\mathbf{x}_{k}\}_{k}$ be the sequence generated by Algorithm 1 with $B\geq L$. Suppose that Assumptions 3.1-3.4 hold. Let $\alpha$ and $\delta$ satisfy (3.13). If $\mathbf{x}_{0}\in\mathcal{N}(\delta)$, there exists constants $M_{2},Q,L_{2}$ such that
$\min_{1\leq k\leq K}\mathbb{E}[\|\mathrm{grad}f(\bar{x}_{k})\|^{2}]$
$\leq$ $\frac{4(f(\bar{x}_{0})-f_{*})}{\alpha K}+4(\mathcal{G}_{1}\alpha^{2}+\mathcal{G}_{2}\alpha^{3})+6(\rho_{3}\alpha^{2}+\rho_{2}\nu^{2}\tau^{2})$
$\frac{48L^{2}}{(1-\sigma_{2}^{2})K}+\frac{12}{n}(1+\rho_{2}\tau^{2})(\nu^{2}\tau+\rho_{3}n\frac{\alpha^{4}}{\tau}),$
where $\rho_{2}:=\frac{6(1+\sigma_{2}^{2})}{(1-\sigma_{2}^{2})^{2}},\rho_{3}:=\rho_{2}(8CL^{2}+4DL^{2})$ and $\mathcal{G}_{1},\mathcal{G}_{2}$ are defined as:
(3.16) $\mathcal{G}_{1}:$ $=\frac{3}{2}CL^{2}+(\frac{3}{2}M_{2}^{2}+\frac{3}{2}(\sqrt{n}L_{2}+8Q)^{2})C^{2}$
$+24Q^{2}D^{2}+\frac{1}{4}LL_{2}^{2}+LD,$
$\mathcal{G}_{2}:$ $=2(8Q+\sqrt{n}L_{2}+M_{2})^{2}LC^{2}+8Q^{2}D^{2}L+2M_{2}C^{2}L.$
The following corollary addresses the finite-time convergence rate of DPRSRM with specific choices of the algorithmic parameters $\alpha$ and $\tau$.
Corollary 3.8.
Let $\{\mathbf{x}_{k}\}_{k}$ be the sequence generated by Algorithm 1 with $B\geq L$. Suppose that Assumptions 3.1-3.4 hold. Let $\alpha=\frac{1}{K^{1/3}}$, $\tau=\frac{1}{K^{2/3}}$ and $\delta\leq\min\{\frac{R}{2},\frac{R(1-\sigma_{2})}{2}\}$. If $\mathbf{x}_{0}\in\mathcal{N}(\delta)$, and $K\geq\max\left\{4L,\sqrt{nD}/\delta,\frac{R\sqrt{nD}}{(R(1-\sigma_{2})-2\delta)\delta}\right\}$, it holds that
(3.17) $\mathbb{E}\left[\frac{1}{n}\|\bar{\mathbf{x}}_{K}-\mathbf{x}_{K}\|^{2}\right]$ $\leq\frac{C}{K^{2/3}},$
$\min_{0\leq k\leq K}\mathbb{E}[\|\mathrm{grad}f(\bar{x}_{k})\|^{2}]$ $\leq\frac{\Gamma_{1}}{K^{2/3}}+\frac{\Gamma_{2}}{K}+\frac{6\rho_{2}\nu^{2}}{K^{4/3}},$
where $\Gamma_{1}=4(f(\bar{x}_{0})-f_{*})+4\mathcal{G}_{1}+6\rho_{3}+\frac{12\sigma}{n}(1+\rho_{2})$ and $\Gamma_{2}=4\mathcal{G}_{2}+\frac{48L^{2}}{(1-\sigma_{2}^{2})}$. As a consequence, DPRSRM obtains an $\epsilon$-stationary point with at most
$\mathcal{K}:=\mathcal{O}(\max\{\mathcal{K}_{1},\mathcal{K}_{2},\mathcal{K}_{3}\})$
iterations. Here, $\mathcal{K}_{1},\mathcal{K}_{2},\mathcal{K}_{3}$ are given as follows:
(3.18) $\mathcal{K}_{1}:$ $=(C+\Gamma_{1})^{1.5}\epsilon^{-\frac{3}{2}},\mathcal{K}_{2}:=\Gamma_{2}\epsilon^{-1},$
$\mathcal{K}_{3}:$ $=(6\rho_{2}\nu^{2})^{3/4}\epsilon^{-\frac{3}{4}}.$
According to Corollary 3.8, DPRSRM achieves an oracle complexity of $\mathcal{O}(\epsilon^{-\frac{3}{2}})$, outperforming existing methods [7, 40], which have an oracle complexity of at most $\mathcal{O}(\epsilon^{-2})$.
4 Outline of convergence analysis
As shown in Section 3.3, the convergence analysis consists of two critical components: the consensus error and the optimality measure, corresponding to Theorem 3.6 and Theorem 3.7, respectively. In this section, we will outline the proof.
4.1 Consensus error
This subsection addresses the consensus error in Theorem 3.6. In Algorithm 1, the update of the main iterate $x_{k+1}$ involves the projection on $\mathcal{M}$. Due to the nonconvex nature of compact submanifolds, the projection is not always unique. Before proving Theorem 3.6, we will first demonstrate that the projection operator in Algorithm 1 is well-defined, meaning that the points being projected are always ensured to be within a neighborhood that belongs to $\bar{U}_{\mathcal{M}}(R)$.
We first investigate the uniform boundedness of $\|\mathrm{\bf s}_{k}\|$ in the following lemma.
Lemma 4.1.
Let $\{\mathbf{x}_{k}\}_{k}$ be the sequence generated by Algorithm 1. Suppose that Assumptions 3.1 and 3.3 hold. Then for all $k$, it holds that
(4.1) $\|\mathrm{\bf s}_{k}\|^{2}\leq nD,~{}~{}D:=\frac{(3L)^{2}}{(1-\sigma_{2})^{2}}.$
The following lemma demonstrates that the iterates ${\mathbf{x}_{k}}$ will always remain in the neighborhood $\mathcal{N}(\delta)$ under certain conditions.
Lemma 4.2.
Suppose that Assumptions 3.1 and 3.3 hold. Let $x_{i,k}$ be generated by Algorithm 1. Let $\alpha$ and $\delta$ satisfy (3.13). If $\mathbf{x}_{0}\in\mathcal{N}(\delta)$, then for any $k\geq 1$, it holds that $\mathbf{x}_{k}\in\mathcal{N}(\delta)$ and
(4.2) $\sum_{j=1}^{n}{W_{ij}}x_{j,k}-\alpha v_{i,k}$ $\in\bar{U}_{\mathcal{M}}(2\delta),~{}i=1,\cdots,n.$
Now we give the proof of Theorem 3.6.
Proof 4.3 (Proof of Theorem 3.6).
Since $\mathbf{x}_{0}\in\mathcal{N}(\delta)$, it follows from Lemma 4.2 that for any $k>0$, $\mathbf{x}_{k}\in\mathcal{N}(\delta)$ and
(4.3) $\sum_{j=1}^{n}{W_{ij}}x_{j,k}-\alpha v_{i,k}\in\bar{U}_{\mathcal{M}}(2\delta),~{}i=1,\cdots,n.$
By the definition of $\bar{\mathbf{x}}_{k+1}$, we have
$\|\mathbf{x}_{k+1}-\bar{\mathbf{x}}_{k+1}\|$ $\leq\|\mathbf{x}_{k+1}-\bar{\mathbf{x}}_{k}\|$
$=\|\mathcal{P}_{\mathcal{M}^{n}}(\mathbf{W}\mathbf{x}_{k}-\alpha\mathbf{v}_{k})-\mathcal{P}_{\mathcal{M}^{n}}(\hat{\mathbf{x}}_{k})\|$
$\leq\frac{R}{R-2\delta}\|\mathbf{W}\mathbf{x}_{k}-\alpha\mathbf{v}_{k}-\hat{\mathbf{x}}_{k}\|$
$\leq\frac{R}{R-2\delta}\sigma_{2}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|+\frac{R}{R-2\delta}\sqrt{nD}\alpha$
$\leq\rho\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|+\frac{\sqrt{nD}\alpha}{\sigma_{2}},$
where the first inequality follows from the optimality of $\bar{\mathbf{x}}_{k+1}$, the second inequality utilizes (4.3) and the $\frac{R}{R-2\delta}$-Lispchitz continuity of $\mathcal{P}_{\mathcal{M}}$ over $\bar{U}_{\mathcal{M}}(2\delta)$. Then
(4.4) $\|\bar{\mathbf{x}}_{k+1}-\mathbf{x}_{k+1}\|$ $\leq\rho\|\bar{\mathbf{x}}_{k}-\mathbf{x}_{k}\|+\frac{\sqrt{nD}\alpha}{\sigma_{2}}$
$\leq\rho^{k+1}\|\bar{\mathbf{x}}_{0}-\mathbf{x}_{0}\|+\frac{\sqrt{nD}}{(1-\rho)\sigma_{2}}\alpha.$
By the initialized strategy of $\mathbf{x}_{0}$ in Algorithm 1, we have $\mathbf{x}_{0}=\bar{\mathbf{x}}_{0}$. The proof is completed.
4.2 Optimality error
In this subsection, we outline the proof of Theorem 3.7. For ease of notation, let us denote
$\hat{g}_{k}$ $:=\frac{1}{n}\sum_{i=1}^{n}\mathrm{grad}f_{i}\left(x_{i,k}\right),\quad\hat{\mathbf{g}}_{k}:=\left(\mathbf{1}_{n}\otimes I_{d}\right)\hat{g}_{k}.$
By applying the Lipschitz-type inequalities on compact submanifolds from Section 2.2 and combining them with the above lemma, we can demonstrate a sufficient decrease in $f$.
Lemma 4.4.
Let $\{\mathbf{x}_{k}\}_{k}$ be the sequence generated by Algorithm 1. Suppose that Assumptions 3.1 and 3.3 hold. Let $\alpha$ and $\delta$ satisfy (3.13). If $\mathbf{x}_{0}\in\mathcal{N}(\delta)$, it follows that
$f(\bar{x}_{k+1})\leq f(\bar{x}_{k})$ $-\frac{\alpha}{4}\|\mathrm{grad}f(\bar{x}_{k})\|+\mathcal{G}_{1}\alpha^{3}+\mathcal{G}_{2}\alpha^{4}$
$+\frac{3\alpha}{2n}(\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}+\|\hat{\mathbf{d}}_{k}-\mathrm{\bf s}_{k}\|^{2}),$
where $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are constant defined by (3.16).
Let us build the following lemma on the relationship between ${\bf s}_{k}$ and $\hat{\bf d}_{k}$.
Lemma 4.5.
Let $\{\mathbf{x}_{k}\}_{k}$ be the sequence generated by Algorithm 1. Suppose that Assumptions 3.1-3.4 hold. It holds that for any $k$,
(4.5) $\sum_{k=0}^{K}\mathbb{E}[\|\hat{\mathbf{d}}_{k}-\mathrm{\bf s}_{k}\|^{2}]\leq\rho_{2}\sum_{k=0}^{K-1}\tau^{2}\mathbb{E}[\|\mathbf{d}_{k}-\mathrm{grad}f(\mathbf{x}_{k})\|^{2}]$
$+\frac{8nL^{2}}{1-\sigma_{2}^{2}}+\rho_{3}n\alpha^{2}K+\rho_{2}\nu^{2}n\tau^{2}K.$
We also have the gradient estimation error bound.
Lemma 4.6.
Let $\{\mathbf{x}_{k}\}_{k}$ be the sequence generated by Algorithm 1 with $B\geq L$. Suppose that Assumptions 3.1-3.4 hold. Then the expected estimation error of the estimator is bounded by
(4.6) $\sum_{k=0}^{K}\mathbb{E}[\|\mathbf{d}_{k}-\mathrm{grad}f(\mathbf{x}_{k})\|^{2}]\leq\frac{n\nu^{2}}{\tau}+2\nu^{2}\tau K+2\rho_{3}n\frac{\alpha^{4}}{\tau}K,$
where $\rho_{3}$ is defined in Lemma 4.5.
With these preparations, we give the proof of Theorem 3.7.
Proof 4.7 (Proof of Theorem 3.7).
Combining Lemma 4.4 and Lemma 4.5 yields that
(4.7) $\sum_{k=0}^{K}\frac{\alpha}{4}\mathbb{E}[\|\mathrm{grad}f(\bar{x}_{k})\|^{2}]\leq f(\bar{x}_{0})+(\mathcal{G}_{1}\alpha^{3}+\mathcal{G}_{2}\alpha^{4})K$
$+\frac{3\alpha}{2n}\sum_{k=0}^{K}\mathbb{E}[\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}+\|\hat{\mathbf{d}}_{k}-\mathrm{\bf s}_{k}\|^{2}]$
$\leq$ $f(\bar{x}_{1})+(\mathcal{G}_{1}\alpha^{3}+\mathcal{G}_{2}\alpha^{4})K+\frac{3\alpha}{2}(\rho_{3}\alpha^{2}+\rho_{2}\nu^{2}\tau^{2})K$
$+\frac{12L^{2}}{1-\sigma_{2}^{2}}\alpha+\frac{3\alpha}{2n}(1+\rho_{2}\tau^{2})\sum_{k=0}^{K}\mathbb{E}[\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}].$
Incorporating Lemma 4.6 into (4.7) completes the proof.
Proof 4.8 (Proof of Corollary 3.8).
Given the choice of $\alpha$ and $K\geq\max\left\{4L,\sqrt{nD}/\delta,\frac{R\sqrt{nD}}{(R(1-\sigma_{2})-2\delta)\delta}\right\}$, we can infer that $\alpha$ satisfies the condition (3.13). Therefore, it follows from Theorem 3.7 that
(4.8) $\min_{1\leq k\leq K}\mathbb{E}[\|\mathrm{grad}f(\bar{x}_{k})\|^{2}]$
$\leq$ $\frac{4(f(\bar{x}_{0})-f_{*})}{\alpha K}+4(\mathcal{G}_{1}\alpha^{2}+\mathcal{G}_{2}\alpha^{3})+6(\rho_{3}\alpha^{2}+\rho_{2}\nu^{2}\tau^{2})$
$\frac{48L^{2}}{(1-\sigma_{2}^{2})K}+\frac{12}{n}(1+\rho_{2}\tau^{2})(\nu^{2}\tau+\rho_{3}n\frac{\alpha^{4}}{\tau})$
$\leq$ $\frac{4(f(\bar{x}_{0})-f_{*})+4\mathcal{G}_{1}+6\rho_{3}+\frac{12\sigma}{n}(1+\rho_{2})}{K^{2/3}}$
$+(4\mathcal{G}_{2}+\frac{48L^{2}}{(1-\sigma_{2}^{2})})\frac{1}{K}+\frac{6\rho_{2}\nu^{2}}{K^{4/3}}.$
The proof is completed.
5 Numerical experiments
In this section, we compare our proposed DPRSRM with DRSGD in [7] and DRPGD in [12] on decentralized principal component analysis. It is important to note that the original DRPGD is a deterministic algorithm that utilizes the local full gradient. To adapt it to the online setting, we replace the full gradient with a stochastic gradient for the local updates. The numerical results on decentralized low-rank matrix completion are provided in the supplementary material.
5.1 Decentralized principal component analysis
The decentralized principal component analysis (PCA) problem can be expressed mathematically as follows:
(5.1) $\min_{\mathbf{x}\in\mathcal{M}^{n}}-\frac{1}{2n}\sum_{i=1}^{n}\text{tr}(x_{i}^{\top}A_{i}^{\top}A_{i}x_{i}),\;\;\text{ s.t. }\;\;x_{1}=\ldots=x_{n},$
where $\mathcal{M}$ is the Stiefel manifold $\text{St}(d,r)$, $A_{i}\in\mathbb{R}^{m_{i}\times d}$ is the data matrix corresponding to the $i$-th node, and $m_{i}$ denotes the number of samples. It is worth noting that if $x^{*}$ is a solution to this problem, then any transformation of $x^{*}$ by an orthogonal matrix $Q\in\mathbb{R}^{r\times r}$ is also a valid solution. The distance between two points $x$ and $x^{*}$ is then calculated as:
$d_{s}(x,x^{*}):=\min_{Q\in\mathbb{R}^{r\times r},\;Q^{\top}Q=QQ^{\top}=I_{r}}\;\|xQ-x^{*}\|.$
5.1.1 Synthetic dataset
In our study, we set the parameters as follows: $m_{1}=\ldots=m_{n}=1000$, $d=10$, and $r=5$. A matrix $B\in\mathbb{R}^{1000n\times d}$ is generated, and its singular value decomposition (SVD) is performed, yielding $B=U\Sigma V^{\top}$, where $U\in\mathbb{R}^{1000n\times d}$ and $V\in\mathbb{R}^{d\times d}$ are orthogonal matrices, and $\Sigma\in\mathbb{R}^{d\times d}$ is a diagonal matrix. To control the distribution of singular values, we define $\tilde{\Sigma}=\text{diag}(\gamma^{j})$ with $\gamma$ chosen from the interval $(0,1)$. The matrix $A$ is then formed as $A=U\tilde{\Sigma}V^{\top}\in\mathbb{R}^{1000n\times d}$. The matrices $A_{i}$ are derived by partitioning the rows of $A$ into $n$ equally sized subsets. It can be shown that the first $r$ columns of $V$ represent the solution to (5.1). In our experiments, the parameters $\gamma$ and $n$ are set to 0.8 and 8, respectively.
We employ fixed step sizes for all algorithms. The step size is set to $\alpha=\frac{\hat{\beta}}{\sqrt{K}}$ with $K$ being the maximal number of iterations. The grid search is utilized to find the best $\hat{\beta}$ for each algorithm. The momentum parameter is chosen as $\tau=0.999$. The batch size in each node is set as $10$ and the maximum iteration is set $K=2000$. The clipping constant is set as $B=10^{8}$. We choose the polar decomposition as the retraction operator for DRSGD. We test several graph matrices to model the topology across the nodes, namely, the Erdos-Renyi (ER) network with probability $p=0.3,0.6$, and the Ring network. Throughout this section, we select the mixing matrix $W$ to be the Metropolis constant edge weight matrix [29].
Figure 1: Numerical results on the synthetic dataset with different network graphs.
Figure 2: Results on the synthetic dataset with ER $p=0.6$.
Firstly, we test all algorithms with different network graphs, namely, Ring, ER $p=0.3$, and ER $p=0.6$. The results are shown in Figure 1. We see that there is not much difference among different graphs except for the consensus error. This is consistent with the existing results for DRSGD [7] and DPRGD [12]. Secondly, Figure 2 presents a comparison among the three algorithms. Our DPRSRM outperforms the other two, with DRSGD and DPRGD showing comparable performance.
Figure 3: Results on Mnist dataset with ER $p=0.3$.
5.1.2 Mnist dataset
We also conduct numerical tests on the Mnist dataset [23]. The training images consist of 60000 handwritten images of size $32\times 32$ and are used to generate $A_{i}$’s. We first normalize the data matrix by dividing 255 and randomly split the data into $n=8$ nodes with equal cardinality. Then, each node holds a local matrix $A_{i}$ of dimension $\frac{60000}{n}\times 784$. We compute the first 5 principal components, i.e., $d=784,r=5$.
For all algorithms, we use the fixed step sizes $\alpha=\frac{\hat{\beta}}{60000}$ with a best-chosen $\hat{\beta}$, batch size 1500 and momemtum parameter $\tau=0.999$. Similar to the synthetic setting, our DPRSRM algorithm demonstrates superior performance compared to other algorithms in terms of objective function values, gradient norms, and distances to the optimal solution. It is important to note that due to the use of stochastic gradients, consensus among the algorithms may be affected by noise, which can be reduced by using a smaller step size.
5.2 Low-rank matrix completion problem
The low-rank matrix completion (LRMC) problem aims to reconstruct a matrix $A\in\mathbb{R}^{d\times T}$ with low rank from its partially observed entries. Let $\Omega$ represent the set of indices corresponding to the observed entries in $A$. The LRMC problem of rank $r$ can be expressed as:
(5.2) $\min_{X\in{\rm Gr}(d,r),V\in\mathbb{R}^{r\times T}}\quad\frac{1}{2}\|\mathcal{P}_{\Omega}(XV-A)\|^{2},$
where $\text{Gr}(d,r)$ represents the Grassmann manifold of $r$-dimensional subspaces in $\mathbb{R}^{d}$, and $\mathcal{P}_{\Omega}$ is a projection operator that selects the elements of $A$ indexed by $\Omega$, setting the rest to zero.
In a decentralized context, assume the matrix $\mathcal{P}_{\Omega}(A)$ is divided into $n$ equal parts by columns, labeled as $A_{1},A_{2},\ldots,A_{n}$, each part corresponding to a different node. Replacing the Grassmann manifold constraint with the Stiefel manifold, the decentralized LRMC problem [12] is therefore formulated as:
(5.3) $\min$ $\frac{1}{2}\sum_{i=1}^{n}\|\mathcal{P}_{\Omega_{i}}(X_{i}V_{i}(X)-A_{i})\|^{2},$
$\mathrm{s.t.}$ $X_{1}=X_{2}=\cdots=X_{n},$
$X_{i}\in{\rm St}(d,r),\quad\forall i\in[n],$
where $\Omega_{i}$ is the corresponding indices set of $\Omega$ and $V_{i}(X):={\rm argmin}_{V}\|\mathcal{P}_{\Omega_{i}}(XV-A_{i})\|$.
For numerical tests, we consider random generated $A$. To be specific, we first generate two random matrices $L\in\mathbb{R}^{d\times r}$ and $R\in\mathbb{R}^{r\times T}$, where each element obeys the standard Gaussian distribution. For the indices set $\Omega$, we generate a random matrix $B$ with each element following from the uniform distribution, then set $\Omega_{ij}=1$ if $B_{ij}\leq\nu$ and 00 otherwise. The parameter $\nu$ is set to $r(d+T-r)/(dT)$. In the implementations, we set $T=1000,d=50,r=10$, and $\alpha=\frac{\hat{\beta}}{\sqrt{K}}$ for all algorithms with $K$ being the maximal number of iterations. $\hat{\beta}$ is tuned to get the best performance for each algorithm individually. The Ring graph is used. The results are reported in Figure 4, where DPRGD is omitted due to its similar performance with DRSGD. We see that DPRSRM outperforms DRSGD.
Figure 4: Numerical results for the decentralized LRMC problem with the Ring graph.
6 Conclusions and Limitations
This work develops a decentralized projection Riemannian stochastic recursive momentum method by assuming that each node has access to a stochastic first-order oracle. Our algorithm leverages local hybrid variance reduction and gradient tracking to achieve a lower oracle complexity compared with the existing online methods. It requires only $\mathcal{O}(1)$ gradient evaluations per iteration for each local node and does not require restarting with a large batch gradient.
Acknowledgements
This effort was supported by the SciAI Center, and funded by the Office of Naval Research (ONR), under Grant Number N00014-23- 1-2729 (J.H.), as well as by the National Natural Science Foundation of China (NSFC) grants 12401419 (K.D.).
Appendix A Technical Lemmas
Lemma A.1.
Given any vectors $a,b\in\mathbb{R}^{n}$, it holds that
(A.1) $\langle a,\frac{b}{\|b\|}\rangle\geq\|a\|-2\|a-b\|.$
Proof A.2.
It follows from the Cauchy inequality that
(A.2) $\langle a,\frac{b}{\|b\|}\rangle$ $=\langle a-b,\frac{b}{\|b\|}\rangle+\|b\|\geq-\|a-b\|+\|b\|\geq-2\|a-b\|+\|a\|,$
where the second inequality use $\|a\|\leq\|a-b\|+\|b\|$.
Lemma A.3 ([36], Lemma 2).
Let $u_{k}$ and $w_{k}$ be two positive scalar sequences such that for all $k\geq 1$
(A.3) $u_{k}\leq\eta u_{k-1}+w_{k-1},$
where $\eta\in(0,1)$ is the decaying factor. Then we have
(A.4) $\sum_{k=0}^{K}u_{k}\leq\frac{u_{0}}{1-\eta}+\frac{1}{1-\eta}\sum_{k=0}^{K-1}w_{k}.$
The following inequality is the control of the distance between the Euclidean mean $\hat{x}$ and the manifold mean $\bar{x}$ by the square of consensus error.
Lemma A.4 ([12]).
For any $\mathbf{x}\in\mathcal{M}^{n}$ satisfying $\|x_{i}-\bar{x}\|\leq{\delta}$, $i\in[n]$, we have
(A.5) $\|\bar{x}-\hat{x}\|\leq M_{2}\frac{\|\mathbf{x}-\bar{\mathbf{x}}\|^{2}}{n},$
where $M_{2}=\max_{x\in{\bar{U}_{\mathcal{M}}({\delta})}}\|D^{2}\mathcal{P}_{\mathcal{M}}(x)\|_{\rm op}$.
The following Lipschitz-type inequality for the projection operator $\mathcal{P}_{\mathcal{M}}(\cdot)$ is crucial in the analysis of projection-based methods.
Lemma A.5 ([12]).
For any $x\in\mathcal{M},u\in\{u\in\mathbb{R}^{d\times r}:\|u\|\leq{\delta}\}$, there exists a constant $Q$ such that
(A.6) $\|\mathcal{P}_{\mathcal{M}}(x+u)-x-\mathcal{P}_{T_{x}\mathcal{M}}(u)\|\leq Q\|u\|^{2}.$
Appendix B Proof of Section 4
B.1 Proof of Section 4.1
Proof B.1 (Proof of Lemma 4.1).
We prove it by induction on both $\|\mathrm{\bf s}_{k}\|$. By the initial strategy and the update rule, we have $\|\mathrm{\bf s}_{0}\|=\|\mathbf{d}_{0}\|\leq\sqrt{n}L<\frac{3\sqrt{n}L}{1-\sigma_{2}}$. Suppose for some $k\geq 0$ that $\|s_{i,k}\|\leq\frac{3\sqrt{n}L}{1-\sigma_{2}}$. Then, we have
$\|\mathrm{\bf s}_{k+1}-\hat{\mathbf{d}}_{k}\|$ $=\|\mathbf{W}\mathrm{\bf s}_{k}-\hat{\mathbf{d}}_{k}+\mathbf{d}_{k+1}-\mathbf{d}_{k}\|$
$=\|\mathbf{W}\mathrm{\bf s}_{k}-\hat{\mathrm{\bf s}}_{k}+\mathbf{d}_{k+1}-\mathbf{d}_{k}\|$
$=\|[(W-J)\otimes I_{d}]\mathrm{\bf s}_{k}+\mathbf{d}_{k+1}-\mathbf{d}_{k}\|$
$\leq\sigma_{2}\|\mathrm{\bf s}_{k}\|+2\sqrt{n}L$
$\leq\frac{3\sqrt{n}L\sigma_{2}}{1-\sigma_{2}}+2\sqrt{n}L.$
Hence,
$\|\mathrm{\bf s}_{k+1}\|\leq\|\mathrm{\bf s}_{k+1}-\hat{\mathbf{d}}_{k}\|+\|\hat{\mathbf{d}}_{k}\|\leq\frac{3\sqrt{n}L\sigma_{2}}{1-\sigma_{2}}+3\sqrt{n}L\leq\frac{3\sqrt{n}L}{1-\sigma_{2}},$
where we use $\|\hat{d}_{k}\|\leq\frac{1}{n}\sum_{i=1}^{n}\|d_{i,k}\|\leq L$ and $\|\hat{\mathbf{d}}_{k}\|\leq\sqrt{n}\|\hat{d}_{k}\|=\sqrt{n}L$. The proof is completed.
Proof B.2 (Proof of Lemma 4.2).
Let us prove it by induction. Assume that $\mathbf{x}_{k}\in\mathcal{N}(\delta)$. Note that for any $i\in[n]$,
$\|\sum_{j=1}^{n}W_{ij}x_{j,k}-\alpha v_{i,k}-\bar{x}_{k}\|\leq$ $\|\sum_{j=1}^{n}W_{ij}(x_{j,k}-\bar{x}_{k})-\alpha v_{i,k}\|$
$\leq$ $\sum_{j=1}^{n}W_{ij}\|x_{j,k}-\bar{x}_{k}\|+\alpha\|v_{i,k}\|$
$\leq$ $\delta+\sqrt{nD}\alpha\leq 2\delta.$
[…]
$\|\mathbf{x}_{k+1}-\bar{\mathbf{x}}_{k+1}\|\leq\|\mathbf{x}_{k+1}-\bar{\mathbf{x}}_{k}\|$
$=\|\mathcal{P}_{\mathcal{M}^{n}}(\mathbf{W}\mathbf{x}_{k}-\alpha\mathbf{v}_{k})-\mathcal{P}_{\mathcal{M}^{n}}(\hat{\mathbf{x}}_{k})\|$
$\leq\frac{R}{R-2\delta}\|\mathbf{W}\mathbf{x}_{k}-\alpha\mathbf{v}_{k}-\hat{\mathbf{x}}_{k}\|$
$\leq\frac{R}{R-2\delta}\|[(W-J)\otimes I_{d}](\mathbf{x}_{k}-\hat{\mathbf{x}}_{k})-\alpha\mathbf{v}_{k}\|$
$\leq\frac{R\sigma_{2}}{R-2\delta}\|\mathbf{x}_{k}-\hat{\mathbf{x}}_{k}\|+\frac{R\alpha}{R-2\delta}\|\mathbf{v}_{k}\|$
$\leq\frac{R\sigma_{2}}{R-2\delta}\delta+\frac{R\alpha}{R-2\delta}\sqrt{nD}.$
Given $\delta$, one can deduce $\|\mathbf{x}_{k+1}-\bar{\mathbf{x}}_{k+1}\|\leq\delta$ when $\alpha$ satisfy that
(B.1) $\alpha\leq\frac{(R(1-\sigma_{2})-2\delta)\delta}{R\sqrt{nD}}.$
Note that the right hand of the above inequality is quadratic with respect to $\delta$. When $0\leq\delta\leq\frac{R(1-\sigma_{2})}{2}$, the right hand is greater than 0. We complete the proof.
B.2 Proof of Section 4.2
For ease of analysis, we first introduce the expression for the consensus problem. we consider the following consensus problem over $\mathcal{M}$:
(B.2) $\min_{\mathbf{x}}\phi(\mathbf{x}):=\frac{1}{4}\sum_{i=1}^{n}\sum_{j=1}^{n}W_{ij}\|x_{i}-x_{j}\|^{2},~{}\text{s.t.}~{}x_{i}\in\mathcal{M},i\in[n],$
The gradient of $\phi(\mathbf{x})$ is $\nabla\phi(\mathbf{x}):=[\nabla\phi_{1}(\mathbf{x})^{\top},\nabla\phi_{2}(\mathbf{x})^{\top},\cdots,\nabla\phi_{n}(\mathbf{x})^{\top}]^{\top}=(I-\mathbf{W})\mathbf{x}$, where $\nabla\phi_{i}(\mathbf{x}):=x_{i}-\sum_{j=1}^{n}W_{ij}x_{j},i\in[n]$. In particular, the update rule (3.10) of $\mathbf{x}_{k+1}$ can be rewritten as
(B.3) $\mathbf{x}_{k+1}=\mathcal{P}_{\mathcal{M}^{n}}(\mathbf{x}_{k}-\nabla\phi(\mathbf{x})-\alpha\mathbf{v}_{k}).$
Proof B.3 (Proof of Lemma 4.4).
It follows from the Riemannian quadratic upper bound of $f$ in Lemma 3.2 and $L_{g}\leq L$ that
(B.4) $f(\bar{x}_{k+1})\leq f(\bar{x}_{k})+\left<\mathrm{grad}f(\bar{x}_{k}),\bar{x}_{k+1}-\bar{x}_{k}\right>+\frac{L}{2}\|\bar{x}_{k+1}-\bar{x}_{k}\|^{2}$
$=$ $f(\bar{x}_{k})-\frac{\alpha}{n}\sum_{i=1}^{n}\left<\mathrm{grad}f(\bar{x}_{k}),s_{i,k}\right>+\left<\mathrm{grad}f(\bar{x}_{k}),\bar{x}_{k+1}-\bar{x}_{k}+\alpha\hat{s}_{k}\right>+\frac{L}{2}\|\bar{x}_{k+1}-\bar{x}_{k}\|^{2}$
$=$ $f(\bar{x}_{k})-\frac{\alpha}{2n}\sum_{i=1}^{n}(\|\mathrm{grad}f(\bar{x}_{k})\|^{2}+\|s_{i,k}\|^{2}-\|s_{i,k}-\mathrm{grad}f(\bar{x}_{k})\|^{2})$
+$\mathrm{grad}f(\bar{x}_{k}),\bar{x}_{k+1}-\bar{x}_{k}+\alpha\hat{s}_{k}\right>+\frac{L}{2}\|\bar{x}_{k+1}-\bar{x}_{k}\|^{2}$
$\leq$ $f(\bar{x}_{k})-\frac{\alpha}{2}\|\mathrm{grad}f(\bar{x}_{k})\|^{2}-\frac{\alpha}{2n}\|\mathrm{\bf s}_{k}\|^{2}+\frac{\alpha}{2n}\|\mathrm{grad}f(\bar{\mathbf{x}}_{k})-\mathrm{\bf s}_{k}\|^{2}$
+$\mathrm{grad}f(\bar{x}_{k}),\bar{x}_{k+1}-\bar{x}_{k}+\alpha\hat{s}_{k}\right>+\frac{L}{2}\|\bar{x}_{k+1}-\bar{x}_{k}\|^{2},$
where the second inequality utilizes Lemma A.1. According to Young’s inequality, we have
(B.5) $\mathrm{grad}f(\bar{x}_{k}),\bar{x}_{k+1}-\bar{x}_{k}+\alpha\hat{s}_{k}\right>$ =$\mathrm{grad}f(\bar{x}_{k}),\bar{x}_{k+1}-\bar{x}_{k}+\alpha\hat{v}_{k}\right>+\alpha\left<\mathrm{grad}f(\bar{x}_{k}),\hat{s}_{k}-\hat{v}_{k}\right>$
$\leq\frac{\alpha}{4}\|\mathrm{grad}f(\bar{x}_{k})\|^{2}+\frac{1}{\alpha}\|\bar{x}_{k+1}-\bar{x}_{k}+\alpha\hat{v}_{k}\|^{2}+\alpha\left<\mathrm{grad}f(\bar{x}_{k}),\hat{s}_{k}-\hat{v}_{k}\right>.$
Combining (B.4) and (B.5) leads to
(B.6) $f(\bar{x}_{k+1})\leq$ $f(\bar{x}_{k})-\frac{\alpha}{4}\|\mathrm{grad}f(\bar{x}_{k})\|^{2}-\frac{\alpha}{2n}\|\mathrm{\bf s}_{k}\|^{2}+\frac{\alpha}{2n}\underbrace{\|\mathrm{grad}f(\bar{\mathbf{x}}_{k})-\mathrm{\bf s}_{k}\|^{2}}_{a_{1}}$
$+\frac{1}{\alpha}\underbrace{\|\bar{x}_{k+1}-\bar{x}_{k}+\alpha\hat{v}_{k}\|^{2}}_{a_{2}}+\frac{L}{2}\underbrace{\|\bar{x}_{k+1}-\bar{x}_{k}\|^{2}.}_{a_{3}}+\alpha\underbrace{\left<\mathrm{grad}f(\bar{x}_{k}),\hat{s}_{k}-\hat{v}_{k}\right>}_{a_{4}}.$
Now, let us bound $a_{1}$, $a_{2}$, $a_{3}$, and $a_{4}$, respectively. Applying Lemma 3.2 yields
$a_{1}$ $\leq 3\left\|\mathrm{grad}f(\bar{\mathbf{x}}_{k})-\hat{\mathbf{g}}_{k}\right\|^{2}+3\|\hat{\mathbf{g}}_{k}-\hat{\mathbf{d}}_{k}\|^{2}+3\|\hat{\mathbf{d}}_{k}-\mathrm{\bf s}_{k}\|^{2}$
$\leq 3n\|\mathrm{grad}f(\bar{x}_{k})-\hat{g}_{k}\|^{2}+3n\|\hat{g}_{k}-\hat{d}_{k}\|^{2}+3\|\hat{\mathbf{d}}_{k}-\mathrm{\bf s}_{k}\|^{2}$
$\leq 3\sum_{i=1}^{n}\|\mathrm{grad}f(\bar{x}_{k})-\mathrm{grad}f(x_{i,k})\|^{2}+3\sum_{i=1}^{n}\|\mathrm{grad}f(x_{i,k})-d_{i,k}\|^{2}+3\|\hat{\mathbf{d}}_{k}-\mathrm{\bf s}_{k}\|^{2}$
$\leq 3L^{2}\|\bar{\mathbf{x}}_{k}-\mathbf{x}_{k}\|^{2}+3\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}+3\|\hat{\mathbf{d}}_{k}-\mathrm{\bf s}_{k}\|^{2}.$
For $a_{2}$, it follows from the triangle inequality that
(B.7) $\|\bar{x}_{k+1}-\bar{x}_{k}+\alpha\hat{v}_{k}\|\leq\|\bar{x}_{k}-\hat{x}_{k}\|+\|\bar{x}_{k+1}-\hat{x}_{k+1}\|+\|\hat{x}_{k+1}-\hat{x}_{k}+\alpha\hat{v}_{k}\|$
(${\leq}$ $\frac{M_{2}}{n}(\|\bar{\mathbf{x}}_{k}-\mathbf{x}_{k}\|^{2}+\|\bar{\mathbf{x}}_{k+1}-\mathbf{x}_{k+1}\|^{2})+\|\hat{x}_{k+1}-\hat{x}_{k}+\alpha\hat{v}_{k}\|.$
Therefore, it follows from [12, Lemma 5] that there exists a constant $L_{2}$ such that
(B.8) $\|\hat{x}_{k+1}-\hat{x}_{k}+\alpha\hat{v}_{k}\|=\|\frac{1}{n}\sum_{i=1}^{n}(x_{i,k+1}-x_{i,k}+\alpha v_{i,k})\|$
$\leq$ $\frac{1}{n}\|\sum_{i=1}^{n}(x_{i,k+1}-x_{i,k}+\alpha v_{i,k}+\mathrm{grad}\phi_{i}(\mathbf{x}_{k}))\|+\frac{1}{n}\|\sum_{i=1}^{n}\mathrm{grad}\phi_{i}(\mathbf{x}_{k})\|$
(${\leq}$ $\frac{Q}{n}\sum_{i=1}^{n}\|\alpha v_{i,k}+\nabla\phi_{i}(\mathbf{x}_{k})\|^{2}+\frac{1}{n}\|\sum_{i=1}^{n}\mathrm{grad}\phi_{i}(\mathbf{x}_{k})\|$
$\leq$ $\frac{2Q\alpha^{2}}{n}\|{\bf v}_{k}\|^{2}+\frac{2Q}{n}\|\nabla\phi(\mathbf{x}_{k})\|^{2}+\frac{L_{2}}{\sqrt{n}}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}$
$\leq$ $\frac{2Q\alpha^{2}}{n}\|{\bf v}_{k}\|^{2}+\frac{(\sqrt{n}L_{2}+8Q)}{n}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}.$
Plugging (B.8) into (B.7) and using the fact that $\|\mathbf{v}_{k}\|\leq\|\mathrm{\bf s}_{k}\|$ gives
$a_{2}$ $\leq\|\bar{x}_{k+1}-\bar{x}_{k}+\alpha\hat{v}_{k}\|^{2}$
$\leq\frac{3M_{2}^{2}+6(\sqrt{n}L_{2}+8Q)^{2}}{n^{2}}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{4}+\frac{3M_{2}^{2}}{n^{2}}\|\bar{\mathbf{x}}_{k+1}-\mathbf{x}_{k+1}\|^{4}+\frac{24Q^{2}\alpha^{4}}{n^{2}}\|{\bf s}_{k}\|^{4}.$
By [12, Lemma 6] and $\|\hat{v}_{k}\|^{2}\leq\frac{1}{n}\|\mathbf{v}_{k}\|^{2}$, we obtain
$a_{3}\leq$ $\frac{4(8Q+\sqrt{n}L_{2}+M_{2})^{2}}{n^{2}}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{4}+\frac{16Q^{2}\alpha^{4}}{n^{2}}\|{\bf v}_{k}\|^{4}+4\alpha^{2}\|\hat{v}_{k}\|^{2}+\frac{4M_{2}^{2}}{n^{2}}\|\mathbf{x}_{k+1}-\bar{\mathbf{x}}_{k+1}\|^{4}$
$\leq$ $\frac{4(8Q+\sqrt{n}L_{2}+M_{2})^{2}}{n^{2}}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{4}+\frac{16Q^{2}\alpha^{4}}{n^{2}}\|{\bf s}_{k}\|^{4}+\frac{4\alpha^{2}}{n}\|\mathrm{\bf s}_{k}\|^{2}+\frac{4M_{2}^{2}}{n^{2}}\|\mathbf{x}_{k+1}-\bar{\mathbf{x}}_{k+1}\|^{4}.$
Now we are going to bound $a_{4}$, since $s_{i,k}-v_{i,k}\in N_{x_{i,k}}\mathcal{M}$, it follows that
(B.9) $a_{4}=$ $\mathrm{grad}f(\bar{x}_{k}),\frac{1}{n}\sum_{i=1}^{n}\alpha(s_{i,k}-v_{i,k})\right>=\frac{\alpha}{n}\sum_{i=1}^{n}\left<\mathrm{grad}f(\bar{x}_{k})-\mathcal{P}_{T_{x_{i,k}}\mathcal{M}}(\mathrm{grad}f(\bar{x}_{k})),s_{i,k}-v_{i,k}\right>$
$\leq$ $\frac{1}{4n}\sum_{i=1}^{n}\|\mathcal{P}_{T_{\bar{x}_{k}}\mathcal{M}}(\mathrm{grad}f(\bar{x}_{k}))-\mathcal{P}_{T_{x_{i,k}}\mathcal{M}}(\mathrm{grad}f(\bar{x}_{k}))\|^{2}+\frac{\alpha^{2}}{n}\sum_{i=1}^{n}\|\mathcal{P}_{N_{x_{i,k}}\mathcal{M}}(s_{i,k})\|^{2}$
$\leq$ $\frac{L_{2}^{2}}{4n}\sum_{i=1}^{n}\|x_{i,k}-\bar{x}_{k}\|^{2}\|\mathrm{grad}f(\bar{x}_{k})\|^{2}+\frac{\alpha^{2}}{n}\|\mathrm{\bf s}_{k}\|^{2}$
$\leq$ $\frac{L^{2}L_{2}^{2}}{4n}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}+\frac{\alpha^{2}L}{n}\|\mathrm{\bf s}_{k}\|^{2}.$
Combining $a_{1},a_{2},a_{3},a_{4}$ with (B.6) and using the fact that $\alpha<1/(4L)$ imply that
$\mathbb{E}f(\bar{x}_{k+1})\leq f(\bar{x}_{k})-\frac{\alpha}{4}\|\mathrm{grad}f(\bar{x}_{k})\|^{2}-\frac{\alpha}{2n}\|\mathrm{\bf s}_{k}\|^{2}+\frac{\alpha}{2n}a_{1}+\frac{1}{\alpha}a_{2}+\frac{L}{2}a_{3}+\alpha a_{4}$
$\leq$ $f(\bar{x}_{k})-\frac{\alpha}{4}\|\mathrm{grad}f(\bar{x}_{k})\|^{2}+\frac{3L^{2}\alpha}{2n}\|\bar{\mathbf{x}}_{k}-\mathbf{x}_{k}\|^{2}+\frac{3\alpha}{2n}\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}+\frac{3\alpha}{2n}\|\hat{\mathbf{d}}_{k}-\mathrm{\bf s}_{k}\|^{2}$
$+\frac{3M_{2}^{2}+6(\sqrt{n}L_{2}+8Q)^{2}}{n^{2}\alpha}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{4}+\frac{3M_{2}^{2}}{n^{2}\alpha}\|\bar{\mathbf{x}}_{k+1}-\mathbf{x}_{k+1}\|^{4}+\frac{24Q^{2}\alpha^{3}}{n^{2}}\|{\bf s}_{k}\|^{4}.$
$+\frac{2(8Q+\sqrt{n}L_{2}+M_{2})^{2}L}{n^{2}}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{4}+\frac{8Q^{2}\alpha^{4}L}{n^{2}}\|{\bf s}_{k}\|^{4}+\frac{2M_{2}^{2}L}{n^{2}}\|\mathbf{x}_{k+1}-\bar{\mathbf{x}}_{k+1}\|^{4}$
$+\frac{L^{2}L_{2}^{2}\alpha}{4n}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}+\frac{\alpha^{3}L}{n}\|\mathrm{\bf s}_{k}\|^{2}$
Since $\frac{1}{n}\|\mathrm{\bf s}_{k}\|^{2}\leq D$ and $\frac{1}{n}\|\bar{\mathbf{x}}_{k}-\mathbf{x}_{k}\|^{2}\leq C\alpha^{2}$ by Lemma 4.1 and Theorem 3.6, it holds that
$f(\bar{x}_{k+1})\leq f(\bar{x}_{k})$ $-\frac{\alpha}{4}\|\mathrm{grad}f(\bar{x}_{k})\|+\mathcal{G}_{1}\alpha^{3}+\mathcal{G}_{2}\alpha^{4}$
$+\frac{3\alpha}{2n}(\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}+\|\hat{\mathbf{d}}_{k}-\mathrm{\bf s}_{k}\|^{2}),$
where $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are two given by
(B.10) $\mathcal{G}_{1}:$ $=\frac{3}{2}CL^{2}+(\frac{3}{2}M_{2}^{2}+\frac{3}{2}(\sqrt{n}L_{2}+8Q)^{2})C^{2}+24Q^{2}D^{2}+\frac{1}{4}LL_{2}^{2}+LD,$
$\mathcal{G}_{2}:$ $=2(8Q+\sqrt{n}L_{2}+M_{2})^{2}LC^{2}+8Q^{2}D^{2}L+2M_{2}C^{2}L.$
The proof is completed.
Proof B.4 (Proof of Lemma 4.5).
We observe that
(B.11) $d_{i,k+1}-d_{i,k}=$ $\mathrm{grad}f_{i}(x_{i,k+1},\xi_{i,k+1})+(1-\tau)(d_{i,k}-\mathrm{grad}f_{i}(x_{i,k},\xi_{i,k+1}))-d_{i,k}$
$=$ $\mathrm{grad}f_{i}(x_{i,k+1},\xi_{i,k+1})-\mathrm{grad}f_{i}(x_{i,k},\xi_{i,k+1})+\tau(\mathrm{grad}f_{i}(x_{i,k})-d_{i,k})$
$+\tau(\mathrm{grad}f_{i}(x_{i,k},\xi_{i,k+1})-\mathrm{grad}f_{i}(x_{i,k}))$
Then we have that
(B.12) $\mathbf{d}_{k+1}-\mathbf{d}_{k}=$ $\mathrm{grad}f(\mathbf{x}_{k+1})-\mathrm{grad}f(\mathbf{x}_{k})+\tau(\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k})$
$+\tau(\mathrm{grad}f(\mathbf{x}_{k},\mathbf{\xi}_{k+1})-\mathrm{grad}f(\mathbf{x}_{k})).$
The Lipschitz continuity of $\mathrm{grad}f(\mathbf{x})$ and Assumption 3.4 yields
(B.13) $\mathbb{E}[\|\mathbf{d}_{k+1}-\mathbf{d}_{k}\|^{2}]\leq 3L^{2}\mathbb{E}[\|\mathbf{x}_{k+1}-\mathbf{x}_{k}\|^{2}]+3\tau^{2}\mathbb{E}[\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}]+3n\tau^{2}\nu^{2}.$
On the other hand, it holds that
$\|\mathbf{x}_{k+1}-\mathbf{x}_{k}\|$ $\leq\|\mathbf{x}_{k+1}-\mathbf{x}_{k}+(I_{nd}-\mathbf{W})\mathbf{x}_{k}+\alpha\mathbf{v}_{k}\|+\|(I_{nd}-\mathbf{W})\mathbf{x}_{k}+\alpha\mathbf{v}_{k}\|$
$=\|\mathcal{P}_{\mathcal{M}^{n}}(\mathbf{W}\mathbf{x}_{k}-\alpha\mathbf{v}_{k})-(\mathbf{W}\mathbf{x}_{k}-\alpha\mathbf{v}_{k})\|+\|(I_{nd}-\mathbf{W})\mathbf{x}_{k}+\alpha\mathbf{v}_{k}\|$
$\leq 2\|(I_{nd}-\mathbf{W})\mathbf{x}_{k}+\alpha\mathbf{v}_{k}\|\leq 4\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|+2\alpha\|\mathbf{v}_{k}\|,$
where the first inequality is from the triangle inequality and the second inequality is due to the definition of $\mathbf{x}_{k+1}$. Combing with (B.13) yields
(B.14) $\mathbb{E}[\|\mathbf{d}_{k+1}-\mathbf{d}_{k}\|^{2}]\leq 3\tau^{2}\mathbb{E}[\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}]+24L^{2}\mathbb{E}[\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}]+12L^{2}\alpha^{2}\mathbb{E}[\|\mathbf{v}_{k}\|^{2}]+3n\tau^{2}\nu^{2}.$
It follows from the definition of $\hat{\bf d}_{k+1}$ that
${\bf s}_{k+1}-\hat{\bf d}_{k+1}$ $=((I_{n}-J)\otimes I_{d}){\bf s}_{k+1}$
$=((I_{n}-J)\otimes I_{d})((W\otimes I_{d}){\bf s}_{k}+{\bf d}_{k+1}-{\bf d}_{k})$
$=((W-J)\otimes I_{d})({\bf s}_{k}-\hat{\mathbf{d}}_{k})+((I_{n}-J)\otimes I_{d})({\bf d}_{k+1}-{\bf d}_{k}).$
Here we use $((W-J)\otimes I_{d})\hat{\mathbf{d}}_{k}=0$. Note that for any constant $\zeta>0$ and two vectors $a,b$, it holds that $\|a+b\|^{2}\leq(1+\zeta)\|a\|^{2}+(1+\frac{1}{\zeta})\|b\|^{2}$. Using the spectral property of $W$ and combining with (B.14) yields
(B.15) $\mathbb{E}[\|{\bf s}_{k+1}-\hat{\bf d}_{k+1}\|^{2}]\leq(1+\zeta)\sigma_{2}^{2}\mathbb{E}[\|{\bf s}_{k}-\hat{\bf d}_{k}\|^{2}]+(1+\frac{1}{\zeta})\mathbb{E}[\|\mathbf{d}_{k+1}-\mathbf{d}_{k}\|^{2}]$
$\leq$ $(1+\zeta)\sigma_{2}^{2}\mathbb{E}[\|{\bf s}_{k}-\hat{\bf d}_{k}\|^{2}]+(1+\frac{1}{\zeta})(3\tau^{2}\mathbb{E}[\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}]+24L^{2}\mathbb{E}[\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}]+12L^{2}\alpha^{2}\mathbb{E}[\|\mathbf{v}_{k}\|^{2}]+3n\tau^{2}\nu^{2})$
Since $\zeta$ is any positive number, we let $\zeta=\frac{1-\sigma_{2}^{2}}{2\sigma_{2}^{2}}$. Incorporate it into (B.15) to obtain
(B.16) $\mathbb{E}[\|{\bf s}_{k+1}-\hat{\bf d}_{k+1}\|^{2}]$
$\leq$ $\frac{1+\sigma_{2}^{2}}{2}\mathbb{E}[\|{\bf s}_{k}-\hat{\bf d}_{k}\|^{2}]+3\frac{1+\sigma_{2}^{2}}{1-\sigma_{2}^{2}}(\tau^{2}\mathbb{E}[\|\mathrm{grad}f(\mathbf{x}_{k})-\mathbf{d}_{k}\|^{2}]+8L^{2}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}+4L^{2}\alpha^{2}\|\mathbf{v}_{k}\|^{2}+n\tau^{2}\nu^{2}).$
Apply Lemma A.3 to obtain
(B.17) $\sum_{k=0}^{K}\mathbb{E}[\|\hat{\mathbf{d}}_{k}-\mathrm{\bf s}_{k}\|^{2}]$
$\leq$ $\frac{2\mathbb{E}[\|\hat{\mathbf{d}}_{0}-\mathrm{\bf s}_{0}\|^{2}]}{1-\sigma_{2}^{2}}+\frac{6(1+\sigma_{2}^{2})}{(1-\sigma_{2}^{2})^{2}}\sum_{k=0}^{K-1}(\tau^{2}\|\mathbf{d}_{k}-\mathrm{grad}f(\mathbf{x}_{k})\|^{2}+8L^{2}\|\mathbf{x}_{k}-\bar{\mathbf{x}}_{k}\|^{2}+4L^{2}\alpha^{2}\|\mathbf{v}_{k}\|^{2}+n\tau^{2}\nu^{2})$
$\leq$ $\frac{8nL^{2}}{1-\sigma_{2}^{2}}+\frac{6(1+\sigma_{2}^{2})}{(1-\sigma_{2}^{2})^{2}}\sum_{k=0}^{K-1}\tau^{2}\|\mathbf{d}_{k}-\mathrm{grad}f(\mathbf{x}_{k})\|^{2}+\frac{6(1+\sigma_{2}^{2})}{(1-\sigma_{2}^{2})^{2}}(8nCL^{2}\alpha^{2}+4nDL^{2}\alpha^{2}+n\tau^{2}\nu^{2})K,$
where the last inequality follows from (4.1) and (3.14), and uses the fact:
(B.18) $\mathbb{E}[\|\hat{\mathbf{d}}_{0}-\mathrm{\bf s}_{0}\|^{2}]$ $=\mathbb{E}[\|\hat{\mathbf{d}}_{0}-\mathrm{grad}f(\mathbf{x}_{0},\xi_{0})\|^{2}]$
$\leq\sum_{i=1}^{n}\mathbb{E}[\|\hat{d}_{0}-\mathrm{grad}f_{i}(x_{i,0},\xi_{i,0})\|^{2}]$
$\leq\sum_{i=1}^{n}\mathbb{E}[\frac{1}{n}\sum_{j=1}^{n}\|\mathrm{grad}f_{j}(x_{j,0}-\mathrm{grad}f_{i}(x_{i,0},\xi_{i,0})\|^{2}]\leq 4nL^{2}.$
We conclude that (4.5) holds.
Proof B.5 (Proof of Lemma 4.6).
Let us denote
$\Delta_{k}:$ $=\mathbf{d}_{k}-\mathrm{grad}f(\mathbf{x}_{k}),$
$\Delta^{nc}_{k}:$ $=\mathbf{q}_{k}-\mathrm{grad}f(\mathbf{x}_{k}).$
It follows from the update rule of $\mathbf{d}_{k}$ that
(B.19) $\Delta^{nc}_{k}$ $=(1-\tau)\Delta_{k-1}+(1-\tau)(\mathrm{grad}f(\mathbf{x}_{k},\mathbf{\xi}_{k})-\mathrm{grad}f(\mathbf{x}_{k-1},\mathbf{\xi}_{k})+\mathrm{grad}f(\mathbf{x}_{k-1})-\mathrm{grad}f(\mathbf{x}_{k}))$
$+\tau(\mathrm{grad}f(\mathbf{x}_{k},\mathbf{\xi}_{k})-\mathrm{grad}f(\mathbf{x}_{k}))$
Now, let us compare $\|\Delta^{nc}_{k}\|$ and $\|\Delta_{k}\|$. If $\|q_{i,k}\|\leq B$, then
(B.20) $d_{i,k}=q_{i,k}\rightarrow\|\Delta^{nc}_{i,k}\|=\|\Delta_{i,k}\|.$
If $\|q_{i,k}\|>B$, $\|d_{i,k}\|=B$. Since $d_{i,k}$ and $q_{i,k}$ are co-linear, $\|q_{i,k}\|-\|d_{i,k}\|=\|q_{i,k}-d_{i.k}\|$. Therefore
(B.21) $2B\|q_{i,k}-d_{i.k}\|\leq(\|q_{i,k}\|+\|d_{i,k}\|)(\|q_{i,k}\|-\|d_{i,k}\|).$
Then we have that
(B.22) $\|\Delta_{i,k}\|^{2}$ $=\|d_{i,k}\|^{2}-2\langle d_{i,k},\mathrm{grad}f_{i}(x_{i,k})\rangle+\|\mathrm{grad}f_{i}(x_{i,k})\|^{2}$
$\leq\|d_{i,k}\|^{2}-2\langle q_{i,k}-d_{i,k},\mathrm{grad}f_{i}(x_{i,k})\rangle+2\langle q_{i,k},\mathrm{grad}f_{i}(x_{i,k})\rangle+\|\mathrm{grad}f_{i}(x_{i,k})\|^{2}$
$\leq\|d_{i,k}\|^{2}+2L\|q_{i,k}-d_{i,k}\|+2\langle q_{i,k},\mathrm{grad}f_{i}(x_{i,k})\rangle+\|\mathrm{grad}f_{i}(x_{i,k})\|^{2}$
$\leq\|d_{i,k}\|^{2}+(\|q_{i,k}\|+\|d_{i,k}\|)(\|q_{i,k}\|-\|d_{i,k}\|)+2\langle q_{i,k},\mathrm{grad}f_{i}(x_{i,k})\rangle+\|\mathrm{grad}f_{i}(x_{i,k})\|^{2}$
$\leq\|q_{i,k}\|^{2}+2\langle q_{i,k},\mathrm{grad}f_{i}(x_{i,k})\rangle+\|\mathrm{grad}f_{i}(x_{i,k})\|^{2}=\|\Delta^{nc}_{i,k}\|^{2},$
where the third inequality uses $L\leq B$. Combining with (B.20), we have that for any $k$,
(B.23) $\|\Delta_{k}\|^{2}\leq\|\Delta^{nc}_{k}\|^{2}.$
Now take expectation and apply Jensen inequality to obtain:
(B.24) $\mathbb{E}\left[\left\|\Delta_{k}\right\|^{2}|\mathcal{F}_{k}\right]\leq\mathbb{E}\left[\left\|\Delta^{nc}_{k}\right\|^{2}|\mathcal{F}_{k}\right]$
$\leq$ $\mathbb{E}\left[\left\|\mathrm{grad}f\left(\mathbf{x}_{k},\mathbf{\xi}_{k}\right)+\left(1-\tau\right)\left(\mathbf{d}_{k-1}-\mathrm{grad}f\left(\mathbf{x}_{k-1},\mathbf{\xi}_{k}\right)\right)-\mathrm{grad}f\left(\mathbf{x}_{k}\right)\right\|^{2}|\mathcal{F}_{k}\right]$
$\leq$ $\mathbb{E}\left[\|\tau(\mathrm{grad}f(\mathbf{x}_{k},\mathbf{\xi}_{k})-\mathrm{grad}f(\mathbf{x}_{k}))+(1-\tau)(\mathrm{grad}f(\mathbf{x}_{k},\mathbf{\xi}_{k})-\mathrm{grad}f(\mathbf{x}_{k-1},\mathbf{\xi}_{k}))\right.$
$+(1-\tau)(\mathrm{grad}f(\mathbf{x}_{k-1})-\mathrm{grad}f(\mathbf{x}_{k}))+\left.\left(1-\tau\right)\left(\mathbf{d}_{k-1}-\mathrm{grad}f\left(\mathbf{x}_{k-1}\right)\right)\|^{2}|\mathcal{F}_{k}\right]$
$\leq$ $\left(1-\tau\right)^{2}\left\|\Delta_{k-1}\right\|^{2}+\mathbb{E}\left[2\tau^{2}\left\|\mathrm{grad}f\left(\mathbf{x}_{k},\mathbf{\xi}_{k}\right)-\mathrm{grad}f\left(\mathbf{x}_{k}\right)\right\|^{2}|\mathcal{F}_{k}\right]$
$+2(1-\tau)^{2}\mathbb{E}\left[\left\|\mathrm{grad}f\left(\mathbf{x}_{k},\mathbf{\xi}_{k}\right)-\mathrm{grad}f\left(\mathbf{x}_{k}\right)-(\mathrm{grad}f\left(\mathbf{x}_{k-1},\mathbf{\xi}_{k}\right)-\mathrm{grad}f\left(\mathbf{x}_{k-1}\right))\right\|^{2}|\mathcal{F}_{k}\right]$
$\leq$ $\left(1-\tau\right)^{2}\left\|\Delta_{k-1}\right\|^{2}+2\tau^{2}\nu^{2}+2(1-\tau)^{2}\mathbb{E}\left[\left\|\mathrm{grad}f\left(\mathbf{x}_{k},\mathbf{\xi}_{k}\right)-\mathrm{grad}f\left(\mathbf{x}_{k-1},\mathbf{\xi}_{k}\right)\right\|^{2}|\mathcal{F}_{k}\right]$
$\leq$ $\left(1-\tau\right)^{2}\left\|\Delta_{k-1}\right\|^{2}+2\tau^{2}\nu^{2}+2(1-\tau)^{2}L^{2}\alpha^{2}\|\mathbf{x}_{k}-\mathbf{x}_{k-1}\|^{2}$
$\leq$ $\left(1-\tau\right)^{2}\left\|\Delta_{k-1}\right\|^{2}+2\tau^{2}\nu^{2}+2(1-\tau)^{2}L^{2}\alpha^{2}(8\|\mathbf{x}_{k-1}-\bar{\mathbf{x}}_{k-1}\|^{2}+4\alpha^{2}\|\mathbf{v}_{k-1}\|^{2}).$
Take full expectation and Apply Lemma A.3 to obtain
(B.25) $\sum_{k=0}^{K}\mathbb{E}[\|\Delta_{k}\|^{2}]\leq$ $\frac{\mathbb{E}[\|\Delta_{0}\|^{2}]}{1-(1-\tau)^{2}}+\frac{2\tau^{2}\nu^{2}}{1-(1-\tau)^{2}}K$
$+\frac{8(1-\tau)^{2}L^{2}\alpha^{2}\sum_{k=0}^{K-1}(2\|\mathbf{x}_{k-1}-\bar{\mathbf{x}}_{k-1}\|^{2}+\alpha^{2}\|\mathbf{v}_{k-1}\|^{2})}{1-(1-\tau)^{2}}$
$\leq$ $\frac{n\nu^{2}}{\tau}+2\nu^{2}\tau K+\frac{8L^{2}\alpha^{2}}{\tau}\sum_{k=0}^{K-1}(2\|\mathbf{x}_{k-1}-\bar{\mathbf{x}}_{k-1}\|^{2}+\alpha^{2}\|\mathbf{v}_{k-1}\|^{2})$
$\leq$ $\frac{n\nu^{2}}{\tau}+2\nu^{2}\tau K+(16C+8D)nL^{2}\frac{\alpha^{4}}{\tau}K,$
where the last inequality follows from (4.1) and (3.14), the second inequality utilizes $1-(1-\tau)^{2}\geq\tau$, $(1-\tau)^{2}\leq 1$ and uses the fact:
(B.26) $\mathbb{E}[\|\Delta_{0}\|^{2}]=\mathbb{E}[\|\mathrm{grad}f(\mathbf{x}_{0},\xi_{0})-\mathrm{grad}f(\mathbf{x}_{0})\|^{2}]\leq\sum_{i=1}^{n}\mathbb{E}[\|\mathrm{grad}f_{i}(x_{i,0},\xi_{i,0})-\mathrm{grad}f_{i}(x_{i,0})\|^{2}]\leq n\nu^{2}.$
The proof is completed.